OpenStudy (jboat):

Find all solution to the equation sin^2x+sinx=0

myininaya (myininaya):

do you know how to factor sin^2(x)+sin(x)?

OpenStudy (jboat):

no i dont

myininaya (myininaya):

do you know sin^2(x) is the same as sin(x)*sin(x) ?

OpenStudy (jboat):

yes

myininaya (myininaya):

do you also know sin(x)=sin(x)* 1 ?

OpenStudy (jboat):

yes

myininaya (myininaya):

\[\sin^2(x)+\sin(x) \\ \sin(x) \cdot \sin(x)+ \sin(x) \cdot 1 \\ \color{red}{\sin(x)} \cdot \sin(x)+\color{red}{\sin(x)} \cdot 1 \] so do you know how to factor this now?

OpenStudy (jboat):

do i divide by sinx?

myininaya (myininaya):

a b+ a c=a(b+c)

myininaya (myininaya):

do you mean divide both sides by sin(x) ?

myininaya (myininaya):

the only bad thing about that is you may be losing out on some zeros

OpenStudy (jboat):

yea thats what i was thinking but i guess its not correct

myininaya (myininaya):

or solutions

OpenStudy (jboat):

so i do the a b+ ac=a(b+c) thing?

myininaya (myininaya):

yep that is how you will factor what I have asked you to factor that is known as the distributive law/property (don't know if was law or property-can't remember :p )

OpenStudy (jboat):

ok so how do i tell whats a b or c?

myininaya (myininaya):

a is the the factor a*b and a*c have in common

myininaya (myininaya):

what do you think a is here? what factor does sin(x)*sin(x) and sin(x)*1 have in common?

OpenStudy (jboat):

sinx

myininaya (myininaya):

right

OpenStudy (jboat):

ok so a = sinx so b is what?

OpenStudy (jboat):

is it also sinx?

myininaya (myininaya):

yep

OpenStudy (jboat):

and c=1?

myininaya (myininaya):

yep

myininaya (myininaya):

it is just multiplication a(b+c) = a*b+a*c

OpenStudy (jboat):

ok let me see if i got this

OpenStudy (jboat):

sinx(sinx+1)=Sinx x Sinx + Sinx?

myininaya (myininaya):

that extra x between the sin(x)'s shouldn't be there but I don't think you meant to write it there yes \[ \sin^2(x)+\sin(x)=\sin(x)(\sin(x)+1)\]

OpenStudy (jboat):

so is that the answer?

myininaya (myininaya):

so \[\sin(x)(sin(x)+1)=0\] this means we have to solve two equations \[\sin(x)=0 \text{ or } \sin(x)+1=0\]

OpenStudy (jboat):

ok then we use the unit circle and find where y is 0 on the unit circle?

myininaya (myininaya):

sounds great that will work for the first equation sin(x)=0

myininaya (myininaya):

you will also need to find when y is -1 on the unit circle to solve sin(x)+1=0

OpenStudy (jboat):

ok so the solutions are 0, 180, and 270?

myininaya (myininaya):

it says find all solutions the solutions you have given here are solutions in [0,360)

OpenStudy (jboat):

so what are the other solutions?

myininaya (myininaya):

use the period of the sine function is 360 degrees this means keep adding 360 to your solutions above and you will still have a solution to your equation

myininaya (myininaya):

you can do by adding 360*k to each of your solutions in [0,360)

myininaya (myininaya):

where k is a integer

OpenStudy (jboat):

so the solutions are 0*k, 180*k, and 360*k. Where k is an integer

myininaya (myininaya):

the solutions you just told me were 0,180, and 270 I have asked you just to add on +360k to all of these to get all solutions

myininaya (myininaya):

for example the first solution you got was 0 just do 0+360k or 360k

OpenStudy (jboat):

ok so solutions are 0+360k, 180+360k, and 270+360k?

myininaya (myininaya):

yep and you can write the first one as 360k since 0+anything=that anything

OpenStudy (jboat):

ok thank you for your help

myininaya (myininaya):

np

OpenStudy (rockinhood):

Please close this question ^^